### v2

# ROVR-Open-Dataset: A Large-Scale Depth Dataset for Autonomous Driving

Abstract
Depth estimation is a fundamental task for 3D scene understanding in autonomous driving, robotics, and augmented reality. Existing depth datasets, such as KITTI, nuScenes, and DDAD, have advanced the field but suffer from limitations in diversity and scalability. As benchmark performance on these datasets approaches saturation, there is an increasing need for a new generation of large-scale, diverse, and cost-efficient datasets to support the era of foundation models and multi-modal learning. We present ROVR, a large-scale, diverse, and cost-efficient depth dataset designed to capture the complexity of real-world driving. ROVR comprises 200K high-resolution frames across highway, rural, and urban scenarios, spanning day/night and adverse weather conditions. A lightweight acquisition pipeline ensures scalable collection, while sparse but statistically sufficient ground truth supports robust training. Benchmarking with state-of-the-art monocular depth models reveals severe cross-dataset generalization failures: models achieving near-ceiling accuracy on KITTI degrade drastically on ROVR, and even when trained on ROVR, current methods fall short of saturation. These results highlight the unique challenges posed by ROVR—scene diversity, dynamic environments, and sparse ground truth—establishing it as a demanding new platform for advancing depth estimation and building models with stronger real-world robustness. Extensive ablation studies provide a more intuitive understanding of our dataset across different scenarios, lighting conditions, and generalized ability. Dataset will be available at https://xiandaguo.net/ROVR-Open-Dataset/
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding authors
I Introduction
Depth estimation is a cornerstone of 3D scene understanding, providing essential geometric information for a wide range of applications, including autonomous driving [1, 2], robotics [3, 4], and augmented reality (AR) [5]. Accurate depth maps enable downstream tasks such as obstacle detection, motion planning, semantic mapping, and human–machine interaction. Over the past decade, the availability of high-quality annotated datasets has been a major driver of progress in this field, enabling the development of increasingly sophisticated learning-based approaches. Notable examples include large-scale driving datasets such as KITTI [6, 7], nuScenes [8], and DDAD [9], which have served as critical benchmarks for monocular and stereo depth estimation. However, despite their impact, each of these datasets presents limitations in one or more aspects, including scene diversity, annotation density, sensor calibration, and scalability. KITTI [10], while offering high-precision LiDAR depth, is restricted to a relatively narrow set of urban and suburban scenarios, and its limited scale constrains the training of high-capacity models. nuScenes [8] broadens the range of environments and weather conditions but suffers from lower-quality ground truth due to LiDAR calibration inaccuracies. DDAD [9] achieves dense and long-range depth measurements using a sophisticated multi-sensor setup, yet its reliance on costly hardware and complex calibration makes large-scale collection economically challenging.
With the rapid evolution of deep neural networks–particularly transformer-based architectures [11, 12] and multi-modal foundation models–there has been a significant surge in demand for large-scale, diverse, and challenging datasets. As the capacity of models grows, their performance on existing benchmarks is approaching saturation, leaving diminishing room for further improvement without fundamentally new data sources. This trend is especially evident in autonomous driving, where domain-specific datasets can no longer keep pace with the generalization requirements of modern models.
Meanwhile, as long as the model capacity increases, the benchmark performance of emerging models based on these datasets is approaching the limitation in the near future. Facing the new era of foundation models and multi-modal LLMs, we aimed to first propose a very large-scale 3D depth dataset series in diverse driving scenarios, and continuously scale up to other spatial understanding challenges for multi-modal foundation models.
We present a new large-scale, diverse, frame-wise continuous, and cost-efficient public dataset for depth estimation in dynamic outdoor environments as the first version of the dataset with 200K video frames, and will scale up to 1 million very soon. Our collection spans a wide variety of scenes and conditions while maintaining low capture cost through a lightweight acquisition pipeline. Compared to existing datasets, our ground-truth depth is sparse but statistically sufficient for training robust depth estimators. The dataset is fully extensible, enabling future scaling in both volume and scene diversity with minimal additional cost.
Beyond providing large-scale training data, our dataset introduces new challenges for the community: the driving diversity is significantly higher and differs from those in KITTI or DDAD, requiring models to adapt to varying sampling distributions, demanding stronger generalization ability; meanwhile, the depth ground truth density is lower than existing benchmarks, which is harder for model to learn. We summarize key comparison against existing datasets in Table I, present scene-wise distributions in Table II, and show representative samples in Figure 2. The capture setup and sensor configuration are illustrated in Figure 1. We further benchmark standard monocular depth estimation baselines on our dataset (Section IV), demonstrating its utility and the remaining gap to close in challenging scenarios. To provide a technical understanding of the proposed ROVR open dataset, we benchmarked several top-ranked KITTI models in Table IV and domain-specific testing in Table VVI. Notably, SOTA models on KITTI are far from saturating depth estimation metrics on ROVR, highlighting its greater difficulty and broader coverage of real-world scenarios. ROVR provides a crucial platform for developing depth estimation models with stronger generalization to complex, dynamic, real-world scenarios.
TABLE I: Comparison of available autonomous driving depth datasets.
Dataset Scenario Frames Resolution Labels
KITTIDepth [10] Driving 86,000 7,000 1226 $\times$ 370
nuScenes [8] Driving 34,149 5,851 1600 $\times$ 900
[…]
II Related Work
Autonomous Driving Depth Datasets. The KITTI benchmark [6] is one of the earliest and most widely used datasets for driving-scene depth estimation, providing stereo imagery and sparse LiDAR ground truth up to $\sim$80 m (85K/7K, $1226\times 370$). It mainly covers urban and highway scenes under fair weather, limiting diversity in environmental conditions and viewpoints. nuScenes [8] extends to a 360 multi-camera setup with 32-beam LiDAR, radar, and GPS/IMU, collected in Boston and Singapore (34K/5.8K, $1600\times 900$). It offers broader variety—including night and rain—but yields sparser per-image depth due to lower-resolution LiDAR. DDAD [9] addresses range and density constraints via long-range (250 m) high-density LiDAR and high-resolution cameras, providing near-dense depth for diverse urban scenes across multiple continents (17K/4.1K, $1936\times 1216$).
Other Depth Datasets. A variety of datasets target depth estimation beyond autonomous driving scenarios, covering indoor, outdoor, and mixed environments, as well as synthetic domains. NYUv2 [13] provides 795 training and 654 test frames at $640\times 480$ resolution for indoor RGB-D scenes. ScanNet [14] offers a large-scale collection of 2.5M frames with dense annotations for mixed indoor environments at the same resolution. MegaDepth [15] supplies $\sim$128K internet-photo-derived frames via multi-view stereo, spanning diverse indoor/outdoor scenes. TUM RGB-D [16] focuses on benchmarking depth and camera pose estimation with $640\times 480$ resolution RGB-D sequences. Synthetic datasets such as SceneNet RGB-D [17] (5M training and 300K test frames, $320\times 240$) enable controlled, noise-free depth generation for pre-training and domain adaptation. Driving-oriented datasets like Cityscapes [18], RobotCar [19], and KITTIDepth [10] expand coverage for outdoor urban scenes, with KITTIDepth offering 85K/7K stereo frames with ground-truth depth. More recent large-scale driving datasets include DrivingStereo [20] (174K/7.7K frames, $1762\times 800$), nuScenes [8] (34K/5.8K frames, $1600\times 900$), DDAD [9] (17K/4.1K frames, $1936\times 1216$), and Waymo [21] (600K frames, $1920\times 1280$). Mixed-scene datasets such as DIODE [22] and DA-2K [23] offer diverse environments with dense depth, while our proposed ROVR dataset provides 193K/10K high-resolution ($1920\times 1280$) driving frames with centimeter-level accuracy. Finally, the Booster dataset [24, 25] covers indoor scenes featuring transparent and reflective objects.
(a) Real vehicle
(b) Sensor suite on vehicle
Figure 1: Illustration of the data collection vehicles: (a)real-world vehicle used for urban field tests; (c) sensor suite on Vehicle 3, including IMU, GNSS/RTK, camera, LiDAR, and computing platform.
Monocular Depth Estimation. The introduction of end-to-end trainable neural networks in MMDE, pioneered by [26], marked a milestone by introducing optimization with the scale-invariant log loss ($\mathrm{SI}_{\log}$). Subsequent advances span convolution-based architectures [27, 28, 29, 30], diffusion-based models [1] and transformer-based models [31, 32, 33, 34, 11], achieving impressive in-domain accuracy—e.g., PackNet [9], DPT [35], and AdaBins [32] report 5–6% relative error on KITTI, approaching LiDAR precision. However, these models often overfit to their training domain and degrade sharply in novel scenes. To improve cross-domain generalization, recent works explore large-scale and diverse training data, as in the Depth-Anything series [36, 23], or adopt generalizable MMDE strategies [37, 38, 39] that incorporate camera awareness [40, 38] or normalize outputs via intrinsics [41, 42, 39]. While geometric pretraining [37] and dataset-specific priors [39] can boost zero-shot accuracy, these methods typically assume noiseless pinhole intrinsics and rely on predefined backprojection, limiting applicability. Benchmarks such as the Monocular Depth Estimation Challenge (MDEC) [43, 44, 45, 46] now track both in-domain and zero-shot performance, with DepthAnything v2 and Marigold [47] as strong recent baselines.
(a) Highway(Normal)
(b) Highway(Night)
(c) Highway(Rainy)
(d) Urban(Normal)
(e) Urban(Night)
(f) Urban(Rainy)
(g) Rural(Normal)
(h) Rural(Rainy)
Figure 2: Data visualization of the ROVR dataset. The first and third rows show RGB images, while the second and fourth rows present the corresponding depth maps projected onto the RGB images.
III ROVR Datasets
In this section, we introduce the ROVR OpenDataset, a large-scale open 3D dataset tailored for autonomous driving, robotics, and 4D perception tasks. We first describe the data acquisition system, where synchronized LiDAR point clouds and high-resolution images are collected. We then outline the sensor configuration and data collection protocol.
III-A Data Acquisition
The dataset was collected using multiple ROVR data collection vehicles deployed across North America, Europe, and Asia. Each vehicle is equipped with an automotive-grade multi-sensor suite capable of reliable operation under diverse real-world conditions, including day/night cycles and various weather scenarios. The system has recorded over 10,000 hours of high-quality driving data covering urban, suburban, and highway environments.
[…]
For accurate spatial alignment, a full calibration is performed between the LiDAR, camera, GNSS, and IMU using standard multi-sensor calibration pipelines. All sensors are hardware-synchronized via GPS-disciplined clocks, achieving a temporal offset of less than 2 ms between modalities. LiDAR data is recorded at 5 Hz, while camera streams operate at 5 Hz, with each frame paired with precise vehicle pose information.
TABLE II: Statistical overview of ROVR Open Datasets.
Train Test
Setting Highway Rural Urban Sum Highway Rural Urban Sum
Normal 114,644 6,411 42,937 163,992 5,533 600 2,080 8,213
Night 20,223 0 1,650 21,873 1,040 0 150 1,190
Rain 1,496 4,041 2,246 7,783 150 299 150 599
Sum 136,393 10,452 46,833 193,648 6,723 899 2,380 10,002
TABLE III: Performance of depth estimation models across KITTI, DDAD, nuScenes, and ROVR. Models trained on existing datasets perform well on their own test sets but fail to generalize to ROVR.
lower is better $\downarrow$ higher is better $\uparrow$
Train Test silog abs_rel log10 rms sq_rel log_rms $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
VA-DepthNet [48]
KITTI [10] KITTI [10] 6.7058 0.0472 0.0205 2.0194 0.1393 0.0739 0.9807 0.9967 0.9991
KITTI [10] ROVR (Ours) 49.6455 0.6345 0.2977 14.9118 11.8008 0.7617 0.0904 0.2190 0.5045
DDAD [49] DDAD [49] 11.7611 0.0940 0.0392 5.0639 0.7429 0.1319 0.9095 0.9820 0.9942
DDAD [49] ROVR (Ours) 37.1334 0.7487 0.1831 13.1557 20.7556 0.5445 0.2309 0.6708 0.8809
nuScenes [8] nuScenes [8] 8.2524 0.0464 0.0181 2.2801 0.3644 0.0873 0.9730 0.9837 0.9910
nuScenes [8] ROVR (Ours) 50.7492 0.4864 0.1641 12.5946 10.4858 0.5217 0.4218 0.7016 0.8429
TABLE IV: Cross-dataset evaluation of three depth estimation models: training on KITTI leads to large performance drops on ROVR, while training on ROVR yields better results.
lower is better $\downarrow$ higher is better $\uparrow$
Train Test silog abs_rel log10 rms sq_rel log_rms $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
VA-DepthNet [48]
KITTI [10] ROVR (Ours) 48.7576 0.5896 0.2360 13.7797 12.2677 0.6443 0.1525 0.4188 0.7517
ROVR (Ours) ROVR (Ours) 29.3324 0.1812 0.0524 6.9541 3.9546 0.2977 0.8862 0.9366 0.9580
DCDepth [50]
KITTI [10] ROVR (Ours) 49.9300 0.6396 0.2895 14.9843 12.4077 0.7433 0.0653 0.2122 0.5887
ROVR (Ours) ROVR (Ours) 27.2769 0.1943 0.0510 6.4558 4.1901 0.2786 0.8876 0.9369 0.9603
IEBins [51]
KITTI [10] ROVR (Ours) 48.5049 0.6078 0.2956 15.2223 10.7613 0.7563 0.0705 0.1990 0.5830
ROVR (Ours) ROVR (Ours) 29.9729 0.2146 0.0664 7.3685 3.7523 0.3106 0.8411 0.9236 0.9520
III-B Data Organization and Alignment
The processed ROVR OpenDataset is released in ROS 2 bag format, with each sequence folder named as <UTC_time>-<device_ID>-<seq_ID>-<code>. Within each sequence, synchronized multi-modal data are stored in modality-specific subdirectories, including $1920\times 1080$ RGB images (images/), LiDAR point clouds in PCD format (pointclouds/), LiDAR-projected depth maps in single-channel PNG (depth/), and perception annotations (annotation/). Auxiliary navigation data include interpolated GNSS/INS poses (ego_poses.json), raw GNSS logs (ego_poses_raw.json), and 100 Hz IMU measurements (imu_data.csv). All files are timestamped at nanosecond resolution for cross-modal synchronization.
Multi-Sensor Calibration.
Each acquisition platform undergoes full intrinsic and extrinsic calibration prior to deployment. Camera intrinsics and distortion coefficients are stored in int.yaml, while LiDAR-to-camera extrinsics are stored in ext.yaml as a Rodrigues rotation vector and translation vector, following a coordinate remapping (x=y,y=z,z=x)(x=-y,\,y=-z,\,z=x). GNSS–IMU lever-arm offsets are estimated from synchronized motion sequences. These parameters enable precise projection of LiDAR points $\mathbf{p}_{l}$ onto the camera image plane:
$\mathbf{p}_{c}\sim\mathbf{K}\left(\mathbf{R}_{l\rightarrow c}\,\mathbf{p}_{l}+\mathbf{t}_{l\rightarrow c}\right),$ (1)
where $\mathbf{K}$ is the camera intrinsic matrix, and $(\mathbf{R}_{l\rightarrow c},\mathbf{t}_{l\rightarrow c})$ are the extrinsic parameters from LiDAR to camera coordinates.
Data Statistics
As summarized in Table II, the ROVR OpenDataset covers a diverse set of driving environments, including highway, rural, and urban scenes under normal, night, and rainy conditions. The training split contains a total of 193,648 frames, while the test split consists of 10,002 frames. Highway driving dominates the data distribution, but the dataset also includes challenging low-light (night) and adverse-weather (rain) scenarios to promote robustness in perception models.
IV Experiments
TABLE V: Depth estimation performance under varying illumination (light conditions). Zero-shot methods are divided into affine-invariant (top) and metric (bottom).
[…]
Domain Model Type Method silog abs_rel log10 rms sq_rel log_rms $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
Normal Marigold v1.1 [47] 59.6965 0.9370 0.2595 14.2185 15.1616 0.7109 0.1993 0.3969 0.6006
DepthAnythingV2 [23] 71.4943 0.9399 0.2721 14.4564 15.3124 0.8032 0.1974 0.3933 0.5955
DepthAnythingV2 [23] 59.5318 0.9390 0.2595 14.2243 15.2125 0.7105 0.1992 0.3968 0.6005
Depth Pro [52] 59.5396 0.9367 0.2592 14.2148 15.1633 0.7098 0.1996 0.3975 0.6013
UniK3D [53] 59.5338 0.9390 0.2595 14.2244 15.2125 0.7105 0.1992 0.3968 0.6005
Zero-Shot UniDepth v2 [54] 59.5325 0.9390 0.2595 14.2243 15.2125 0.7105 0.1992 0.3968 0.6005
VA-DepthNet [48] 29.8942 0.1833 0.0528 7.0396 4.0945 0.3035 0.8884 0.9366 0.9573
DCDepth [50] 29.6740 0.2033 0.0551 7.0320 4.3746 0.3016 0.8756 0.9335 0.9569
Single-Domain IEBins [51] 30.3786 0.2101 0.0648 7.4493 3.7839 0.3125 0.8501 0.9238 0.9509
Night Marigold v1.1 [47] 54.2200 0.9902 0.2771 13.6023 14.9837 0.7303 0.1520 0.3255 0.5256
DepthAnythingV2 [23] 58.0942 0.9896 0.2804 13.6943 15.0012 0.7542 0.1517 0.3248 0.5245
DepthAnythingV2 [23] 54.2149 0.9902 0.2771 13.6024 14.9842 0.7302 0.1520 0.3255 0.5256
Depth Pro [52] 54.2161 0.9902 0.2771 13.6025 14.9838 0.7302 0.1520 0.3255 0.5256
UniK3D [53] 54.2150 0.9902 0.2771 13.6024 14.9842 0.7302 0.1520 0.3255 0.5256
Zero-Shot UniDepth v2 [54] 54.2149 0.9902 0.2771 13.6024 14.9842 0.7302 0.1520 0.3255 0.5256
VA-DepthNet [48] 27.3737 0.1716 0.0493 6.6754 3.4274 0.2770 0.8804 0.9368 0.9617
DCDepth [50] 27.2400 0.1785 0.0492 6.5566 3.5686 0.2755 0.8717 0.9356 0.9636
Single-Domain IEBins [51] 28.3425 0.2081 0.0589 6.8354 3.7938 0.2914 0.8497 0.9263 0.9592
Rainy Marigold v1.1 [47] 57.9063 1.2563 0.3158 14.6217 20.9788 0.8416 0.1607 0.3243 0.4829
DepthAnythingV2 [23] 69.7717 1.2578 0.3286 14.7804 21.0635 0.9267 0.1586 0.3202 0.4772
DepthAnythingV2 [23] 57.8361 1.2563 0.3157 14.6209 20.9781 0.8412 0.1607 0.3243 0.4830
Depth Pro [52] 57.8361 1.2563 0.3157 14.6209 20.9781 0.8412 0.1607 0.3243 0.4830
UniK3D [53] 57.8363 1.2563 0.3157 14.6209 20.9781 0.8412 0.1607 0.3243 0.4830
Zero-Shot UniDepth v2 [54] 57.8361 1.2563 0.3157 14.6209 20.9781 0.8412 0.1607 0.3243 0.4830
VA-DepthNet [48] 25.5215 0.1709 0.0538 6.3357 3.0841 0.2597 0.8684 0.9360 0.9609
DCDepth [50] 25.2770 0.1773 0.0549 6.2672 3.0250 0.2572 0.8597 0.9316 0.9609
Single-Domain IEBins [51] 27.6484 0.2891 0.1025 7.3199 3.2361 0.3233 0.7015 0.9163 0.9532
We conduct experiments on the ROVR Open Dataset to establish baseline performance and to analyze the influence of different factors. The evaluation consists of (i) benchmark results on the full test split to provide a single reference point for future studies, and (ii) domain-specific results under varying illumination and scene types to illustrate different results.
IV-A Implementation Details
We evaluate two families of depth estimation baselines.
Zero-Shot Models. Networks belonging to this category have been trained on web-scale data, either to predict affine-invariant or metric depth, to excel at zero-shot generalization on unseen domains. We select two affine-invariant and four metric baselines: specifically, Marigold v1.1 [47] and DepthAnythingV2 [23] for the former group, KITTI fine-tuned DepthAnythingV2 [23], Depth Pro [52], UniK3D [53], and UniDepth v2 [54] for the latter. We directly use the weights provided by the authors, selecting those yielding the most accurate results for each model.
Single-domain Models. Networks belonging to this category are usually trained to estimate metric depth on a single large-scale dataset, aiming at excelling in a specific domain. We select three representative baselines: VA-DepthNet [48], DCDepth [50], and IEBins [51]. To ensure comparability, we reproduce the variants that adopt the Swin Transformer–Large (Swin-L) [55] backbone and initialize training solely from its pretrained weights. All images are center-cropped to $960\times 544$ for both training and evaluation, and each model is trained for $5$ epochs. Unless otherwise noted, we follow the official training configurations of each method (e.g., learning rate, weight decay). Ground-truth depth is measured in meters; during training and evaluation, depths are clipped to $[1,80]$ meters and invalid pixels are masked. All experiments are conducted on $8\times$ NVIDIA L20 GPUs.
[…]
IV-B Evaluation Metrics
We follow standard monocular depth estimation [11, 1] practice and report six error metrics where lower is better: scale-invariant log RMSE (silog), absolute relative error (abs_rel), log10 error (log10), root mean squared error (rms), squared relative error (sq_rel), and log RMSE (log_rms). We also report three accuracy metrics where higher is better: $\delta_{1}$, $\delta_{2}$, and $\delta_{3}$, representing the fraction of pixels where the prediction is within a factor of $1.25$, $1.25^{2}$, and $1.25^{3}$ of the ground truth.
TABLE VI: Depth estimation performance across scene types. Zero-shot methods are divided into affine-invariant (top) and metric (bottom).
[…]
Domain Model Type Method silog abs_rel log10 rms sq_rel log_rms $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
Highway Marigold v1.1 [47] 59.2410 0.9960 0.2719 14.2594 16.0318 0.7353 0.1784 0.3640 0.5659
DepthAnythingV2 [23] 69.1255 0.9982 0.2821 14.4595 16.1547 0.8102 0.1771 0.3614 0.5620
DepthAnythingV2 [23] 59.0573 0.9983 0.2718 14.2660 16.0903 0.7348 0.1784 0.3640 0.5659
Depth Pro [52] 59.0670 0.9955 0.2715 14.2549 16.0308 0.7341 0.1787 0.3648 0.5668
UniK3D [53] 59.0593 0.9983 0.2718 14.2662 16.0903 0.7348 0.1784 0.3640 0.5659
Zero-Shot UniDepth v2 [54] 59.0581 0.9983 0.2718 14.2661 16.0903 0.7348 0.1784 0.3640 0.5659
VA-DepthNet [48] 26.2820 0.1476 0.0431 6.2185 3.1248 0.2661 0.9106 0.9484 0.9663
DCDepth [50] 26.3602 0.1614 0.0450 6.2690 3.3054 0.2670 0.9015 0.9454 0.9655
Single-Domain IEBins [51] 27.0938 0.1809 0.0565 6.6169 3.0320 0.2807 0.8690 0.9373 0.9613
Rural Marigold v1.1 [47] 56.1643 0.9741 0.2653 13.8753 15.6193 0.7193 0.2016 0.4003 0.5931
DepthAnythingV2 [23] 69.0038 0.9761 0.2786 14.1191 15.7402 0.8198 0.1992 0.3959 0.5872
DepthAnythingV2 [23] 56.1199 0.9741 0.2653 13.8749 15.6189 0.7190 0.2016 0.4003 0.5931
Depth Pro [52] 56.1199 0.9741 0.2653 13.8749 15.6189 0.7190 0.2016 0.4003 0.5931
UniK3D [53] 56.1207 0.9741 0.2653 13.8749 15.6189 0.7190 0.2016 0.4003 0.5931
Zero-Shot UniDepth v2 [54] 56.1199 0.9741 0.2653 13.8749 15.6189 0.7190 0.2016 0.4003 0.5931
VA-DepthNet [48] 29.1695 0.1803 0.0489 6.9052 4.1370 0.2948 0.8980 0.9420 0.9596
DCDepth [50] 29.2167 0.2031 0.0505 6.7946 4.6425 0.2948 0.8851 0.9402 0.9611
Single-Domain IEBins [51] 29.9301 0.2372 0.0731 7.4677 3.8916 0.3211 0.8352 0.9312 0.9550
Urban DepthAnythingV2 [23] 71.9928 0.8663 0.2598 14.2754 14.0632 0.7837 0.2214 0.4299 0.6280
DepthAnythingV2 [23] 59.0757 0.8636 0.2454 14.0271 13.9164 0.6812 0.2240 0.4345 0.6341
Depth Pro [52] 59.0758 0.8635 0.2454 14.0259 13.9145 0.6812 0.2240 0.4346 0.6341
Marigold v1.1 [47] 59.1288 0.8632 0.2454 14.0258 13.9058 0.6815 0.2240 0.4344 0.6340
UniK3D [53] 59.0768 0.8636 0.2454 14.0272 13.9164 0.6812 0.2240 0.4344 0.6341
Zero-Shot UniDepth v2 [54] 59.0760 0.8636 0.2454 14.0271 13.9164 0.6812 0.2240 0.4344 0.6341
VA-DepthNet [48] 38.0108 0.2763 0.0802 9.0505 6.2298 0.3883 0.8131 0.9013 0.9341
DCDepth [50] 36.8839 0.3027 0.0821 8.8468 6.5511 0.3778 0.7930 0.8977 0.9353
Single-Domain IEBins [51] 38.1213 0.3012 0.0916 9.4542 5.7344 0.3911 0.7645 0.8823 0.9246
IV-C Main Results
As shown in Tables III, models trained on KITTI [10], DDAD [49], or nuScenes [8] achieve strong results within their respective domains but suffer severe degradation when evaluated on ROVR, revealing poor cross-dataset generalization. In particular, metrics such as abs_rel increase by an order of magnitude, and accuracy drops drastically (e.g., $\delta_{1}$ falls from >0.98>0.98 on KITTI to <0.25<0.25 on ROVR).
These results expose two key issues. First, existing benchmarks are already saturated—state-of-the-art methods achieve near-ceiling scores on KITTI—yet the same models fail to generalize to ROVR, demonstrating limited robustness. Second, even when trained directly on ROVR, current methods perform far from saturating the benchmark (best $\text{abs\_rel}\sim 0.18$–0.21), underscoring that ROVR presents a substantially more challenging setting. This establishes ROVR as a more realistic and difficult benchmark, motivating the need for new depth estimation approaches capable of handling its scale, diversity, and sparse ground truth.
IV-D Ablation Study
Table V and Table VI present depth estimation performance across different illumination conditions and scene types. We report the results of three typical depth estimation methods to provide a more complete understanding.
Results under varying illumination conditions. Table V shows results under varying illumination conditions—Normal, Night, and Rainy. As expected, all methods perform better in normal lighting conditions, where visibility and depth prediction quality are optimal. However, the performance significantly declines in low-light conditions, especially at night, as seen in the reduction of metrics like $\delta_{1}$ and $\delta_{2}$. VA-DepthNet consistently outperforms other methods in night-time settings, achieving the lowest abs_rel and rms values. The Rainy condition shows the largest performance drop, with all methods facing challenges due to degraded RGB quality and increased LiDAR sparsity. In rainy conditions, IEBins struggles the most, with substantially higher log10 and sq_rel values, indicating poor performance in handling noise and sparse data.
Results across different scene types. Table VI reports results across different scene types—Highway, Rural, and Urban. Highway scenes, characterized by relatively clear and structured environments, yield the best results for all methods, with VA-DepthNet leading the pack in terms of both $\delta_{1}$ and silog scores. This can be attributed to the simpler geometry and better lighting consistency typically found on highways. Conversely, Urban scenes present the greatest difficulty for depth estimation. With complex urban layouts featuring dense clutter, frequent occlusions, and multi-scale objects, IEBins shows the largest performance gap, with noticeably higher error values across all metrics. These results highlight the challenges posed by dynamic and heterogeneous urban environments, where traditional methods struggle to maintain accuracy.
Figure 3: Abs_rel and $\delta_{1}$ with different test identities. The legends illustrate how performance shifts as the test data size increases from 2K to 10K.
Influence of the Test Scale. Fig. 3 demonstrates that as the number of test identities increases from 2K to 10K, the performance curves remain relatively stable. This indicates that the ROVR dataset provides a consistent and unbiased evaluation environment, where scaling the test set does not distort the overall difficulty. Such stability highlights the dataset’s reliability as a robust benchmark for assessing depth estimation methods at a large scale.
Highway (Night) Refer to caption (a) Refer to caption (b) Refer to caption (c) Refer to caption (d) Refer to caption (e)
Highway (Normal) Refer to caption (f) Refer to caption (g) Refer to caption (h) Refer to caption (i) Refer to caption (j)
Highway (Rainy) Refer to caption (k) Refer to caption (l) Refer to caption (m) Refer to caption (n) Refer to caption (o)
Rural (Normal) Refer to caption (p) Refer to caption (q) Refer to caption (r) Refer to caption (s) Refer to caption (t)
Rural (Rainy) Refer to caption (u) Refer to caption (v) Refer to caption (w) Refer to caption (x) Refer to caption (y)
Urban (Night) Refer to caption (z) Refer to caption ($a$) Refer to caption (ab) Refer to caption (ac) Refer to caption (ad)
Urban (Normal) Refer to caption (ae) Refer to caption (af) Refer to caption (ag) Refer to caption (ah) Refer to caption (ai)
Urban (Rainy) Refer to caption (aj) Image Refer to caption (ak) GT Refer to caption (al) VA-Depthnet Refer to caption (am) IEBins Refer to caption (an) DCDepth
Figure 4: Qualitative comparisons of depth estimation results across different scenarios (highway, rural, urban) and weather conditions (night, normal, rainy).
IV-E Qualitative Results
Fig. 4 showcases qualitative samples from the ROVR dataset across multiple scene types (highway, rural, urban) and varying illumination and weather conditions (normal, night, rainy). Examples are drawn from different illumination conditions and scene types to match the settings in Table V and Table VI. These examples highlight the dataset’s challenging nature: adverse weather and nighttime conditions significantly degrade image quality and reduce ground-truth density, while urban environments introduce dense clutter, occlusions, and multi-scale structures. Such diversity and difficulty expose clear failure modes in existing models, indicating that ROVR not only reflects real-world driving challenges but also sets a demanding benchmark for the next generation of depth estimation methods.
V Conclusion
We introduced ROVR, a large-scale, diverse, and cost-efficient depth dataset designed to reflect the complexity of real-world driving. Covering varied scenes, weather, and illumination conditions, ROVR provides sparse yet statistically sufficient ground truth, enabling robust training while imposing greater generalization challenges. Benchmarking strong KITTI SOTA models revealed substantial performance drops, underscoring the dataset’s difficulty and the limitations of existing benchmarks. By offering a challenging and extensible platform, ROVR aims to drive the development of depth estimation methods capable of handling complex, dynamic, and diverse real-world environments.
Acknowledgements. This work was supported by ROVR Labs, Inc.